# A First Look at Information Highlighting in Stack Overflow Answers

### v1

@@ 4.4 RQ4: Can we recommend the highlighted content in SO answers automatically?
Most likely that the website you are trying to scrape is a dynamic website, meaning that it is Javascript generated code and can’t be scraped with only requests and beautifulsoup .
Most likely that the website you are trying to scrape is a dynamic website, meaning that it is Javascript generated code and can’t be scraped with only requests and beautifulsoup .
Missing Identification: Missed the Code tag.
@@
Our models achieve a precision ranging from 0.62 to 0.82 for different formatting types. It is easier to build a model to recommend Code than other types. Models for text formatting types (i.e., Heading, Bold, and Italic) suffer low recall. Our analysis of failure cases indicates that the majority of the failure cases are due to missing identification. One explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge).
@@ 5.1 Implication of our findings
Future search should consider the highlighted content for the downstream tasks that leverage information from the SO answers. In RQ3, we observe that users tend to highlight important information for a provided answer using Text formatting, such as a reminder or warn of in which context or condition the provided solution works or does not work (i.e., Caveat), an update and new edit in answers (i.e., Update). Such content is informative and important to users when learning and applying knowledge from the answers. Therefore, it is substantial to consider such information for the downstream tasks that leverage information extracted from SO answers, such as answer summarization Ren et al. (2019), and API documentation enrichment Li et al. (2018a); Treude and Robillard (2016).
Our findings provide insights to improve future research in developing approaches for automatic information highlight. In RQ4, we investigate the potential of deep learning NER models in recommending information to highlight on SO posts. Code is easier to identify than other types of formatting. Furthermore, all text-related models (i.e., Italic, Bold, and Heading) suffer from low recall. We reveal one reason is long-tail knowledge. The model tends to memorize frequently highlighted words and miss the recognition of the less-frequent (e.g., tail knowledge). Future research is encouraged to leverage techniques Zhang et al. (2021b), e.g., data augmentation and class rebalancing for tail knowledge.
@@ 5.2 Threats to validity
Internal Validity: Our study involved qualitative analysis in RQs. To reduce the bias, each instance was labeled by two of the authors, and discrepancies were discussed until a consensus was reached. We also showed that the level of inter-rater agreement of the qualitative studies is high. In RQ4, we evaluate our models using precision, recall, and F-1 score with partial match. Although, there might be other metrics that could be used for our task., those metrics are commonly used in NER models, and we follow previous studies Esuli and Sebastiani (2010); Jiang et al. (2016). In RQ4, we utilize a NER architecture from spaCy and train it from scratch based on our own dataset. It was compared with BERT, a transformer model for understanding the quality of the performances of separate architectures. We encourage future research to use more advanced pre-trained large language models.
@@ 6 Conclusion
This paper is the first large-scale study of information highlighting in the text description of SO answers. We find that information highlighting is prevalent, with 47.6% of the 31,169,42 studied answers having text highlighted. We propose a terminology to categorize highlighted text in SO answers and find that source code content (e.g., identifiers and programming keywords) are frequently highlighted using Code and other highlighting formats (i.e., Bold and Italic). Besides code, users also tend to highlight updates (e.g., updates of answers), caveats (i.e., a reminder or warning of in which context or condition the provided solution works or does not work), and references. Further studies could put more effort into investigating how to use the highlighted content for downstream tasks (e.g., answer summarizing) that leverage information from SO answers, and provide tools that can suggest highlighted text for SO users. In addition, we also investigate the potential of recommending the content to be highlighted automatically by adopting named entity recognition (NER) models. Our experimental results highlight that our CNN models consistently achieve precision scores between 0.71 and 0.82 across different formatting types, although they exhibit lower recall rates. On the other hand, while BERT demonstrates superior precision, it struggles with even lower recall. It’s worth noting that spaCy’s Code model performs exceptionally well, with an impressive F1 score of 0.71, outperforming other models in our study. Although BERT usually has better performance, in our case the spaCy model based on CNN architecture worked better. This might be attributed to spaCy’s ability to learn specific patterns within our custom dataset that BERT’s transformer model may have missed. In future research, we plan to enhance the transformer models through fine-tuning and work on improving the recall rates for both models.

### v3

@@ 5.4 RQ4: Can we recommend the highlighted content in SO answers automatically?
8 Code Most likely that the website you are trying to scrape is a dynamic website, meaning that it is Javascript generated code and can’t be scraped with only requests and beautifulsoup . Most likely that the website you are trying to scrape is a dynamic website, meaning that it is Javascript generated code and can’t be scraped with only requests and beautifulsoup . Missing Identification: Missed the Code tag.
@@
Our models achieve a precision ranging from 0.50 to 0.72 for different formatting types. It is easier to build a model to recommend Code than other types. Models for text formatting types (i.e., Heading, Bold, and Italic) suffer low recall. Our analysis of failure cases indicates that the majority of the failure cases are due to missing identification. One explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge).
@@ 6 Discussion
6.1 Usefulness of recommendation models
SO encourages the community to improve the quality of posts by editing them and proper edits will be approved by trusted community members[28]. To demonstrate the usefulness of the proposed recommendation models, we randomly sampled 20 new SO posts that are outside of our dataset, and applied our CNN-based model on them to recommend the information to highlight. We found that in 9 posts, the content recommended by our models was already highlighted. For the rest 11 posts, we highlight the content recommended by the model. All our recommended edits in those 11 posts were accepted by the community. For instance, in Figure 7, our model recommends highlighting the subsection title (e.g., Introduction and The Problem) in Bold and highlighting the code with Code. We posted the edits and all of our edits were approved by the community [^10] https://stackoverflow.com/posts/76886399/revisions.
Figure 7: An example of a post where our formatting improvements were approved by the community.
@@ 6.2 Implication of our findings
Future search should consider the highlighted content for the downstream tasks that leverage information from the SO answers. In RQ3, we observe that users tend to highlight important information for a provided answer using Text formatting, such as a reminder or warn of in which context or condition the provided solution works or does not work (i.e., Caveat), an update and new edit in answers (i.e., Update). Such content is informative and important to users when learning and applying knowledge from the answers. Therefore, it is substantial to consider such information for the downstream tasks that leverage information extracted from SO answers, such as answer summarization [36], and API documentation enrichment [20, 44].
Our findings provide insights to improve future research in developing approaches for automatic information highlight. In RQ4, we investigated the potential of deep learning NER models in recommending information to highlight on SO posts. Code is easier to identify than other types of formatting. Furthermore, all text-related models (i.e., Italic, Bold, and Heading) suffer from low recall. We reveal one reason is long-tail knowledge. The model tends to memorize frequently highlighted words and miss the recognition of the less-frequent (e.g., tail knowledge). Future research is encouraged to leverage techniques [57], e.g., data augmentation and class rebalancing for tail knowledge.
@@ 6.3 Threats to validity
Internal Validity: Our study involved qualitative analysis in RQs. To reduce the bias, each instance was labeled by two of the authors, and discrepancies were discussed until a consensus was reached. We also showed that the level of inter-rater agreement of the qualitative studies is high. In RQ4, we evaluate our models using precision, recall, and F-1 score with partial match. Although, there might be other metrics that could be used for our task, those metrics are commonly used in NER models, and we follow previous studies [10, 17]. In RQ4, we investigated the potential of a CNN-based model and trained it from scratch based on our own dataset. We also investigated a pre-trained model BERT and fine-tuned it for our task. However, our finding might not be generalized to other models. We encourage future research to investigate more models for the task.
@@
Content Validity: Content validity refers to the extent to which a measure represents all facets of a given construct. In this study, we focus on Bold, Italic, Heading, Delete, and Code. Stack Overflow also provides other types of formatting to highlight information, such as lists and links. To mitigate the threat, we focus on the most commonly used ones. Future research is encouraged to study more types of formatting.
Construct Validity: One threat to the construct validity is related to how we collected instances for training recommendation models for different formatting types. Although we made diligent efforts to clean up the dataset and eliminate inconsistencies from both code and text data, as discussed in Section 4.4, it is important to acknowledge that it may not have been possible to remove all inconsistencies. This could potentially pose a threat to the validity of our results. However, our failure cases analysis in Section 5.4 reveals that the proportion of misidentification is low, consistently less than 15.17% across all models. This suggests that following the data cleaning process, the models are not easily confused among different types of formatting. This provides evidence regarding the effectiveness of our data cleaning efforts and increases the reliability of our results. Another threat to the construct validity relates to the suitability of our evaluation metrics: Precision, Recall, and F1-measure based on partial match. Other metrics could be used for evaluating the recommendation system. We selected those metrics because these metrics have been used in many NER studies [10, 17]. This threat is limited by the research.
7 Related work
7.1 Information highlighting
Previous research has explored the benefits of highlighting information in various domains over the last decades [27, 34, 42, 49, 50]. Wu and Yuan have shown that highlighting could reduce the cognitive load and thus reduce reading time [50]. Jorge et al. investigated the impact of highlighting text in the text classification tasks in the machine learning area and found that highlighting is effective in reducing classification effort for humans [34]. Similarly, Nguyen et al. developed a tool to explain the output ML models by highlighting the portion of text and demonstrated its effectiveness in model explanation [27]. Various studies have been done to investigate the impact of information highlighting in software engineering, such as code comprehension [4, 14, 30, 37], source code evolution [9]. For instance, Advait Sarkar investigated the effect of syntax highlighting on program comprehension and its interaction with programming experience [37], and found that syntax highlighting significantly improves task completion time. Different from prior studies that focus on understanding the influence of information highlighting, we focus on understanding the practice of information highlighting on SO.
7.2 Knowledge retrieval on Stack Overflow
Several studies have investigated developers’ challenges and solutions in retrieving knowledge from technical Q&A sites and provided some tools to facilitate the process [12, 26, 51, 56]. For instance, Gottipati et al. found that in software forums it is a painstaking process for users to manually search through answers in various threads of posts [12], and they developed a customized search engine to find relevant answers from various software forums. Similarly, Xu et al. developed a technique called AnswerBot, which takes input as the technical question and generates an answer summary for the question [51]. Zhang et al. pointed out an issue of current comment ranking and display mechanism on SO (e.g., a large portion of useful comments are not visible to users) and proposed an alternative ranking mechanism to alleviate the issue [56]. Sarah and Treude investigated the possibility of using two existing techniques, wordpartten and lexrank, to identify the essential sentences from a long SO answer [26]. Different from prior studies that focus on retrieving relevant or important answers from SO, we investigate what and how information is highlighted on SO. Our study could provide insights to enhance existing techniques.
7.3 Recommendation system for Stack Overflow
Previous studies developed different recommendation tools to ease the use of the Q&A sites, such as recommending hyperlinks [21, 22], tags [15, 23, 35, 46, 48], and similar questions [47, 54]. Li et al. developed LinkLiv to recommend hyperlinks [22], which uses multiple features, including hyperlink co-occurrences in Q&A discussions, and locations (e.g., question, answer, or comment). Wang et al. developed EnTagRec, which combines two complementary lines in the statistics community Bayesian and frequentist [48]. Wang et al. developed SOTagRec combining a convolutional neural network and collaborative filtering model to infer tags for new postings by studying the historical postings and their tags [46]. Maity et al. developed DeepTagRec [23], which learns from the content representation of question title and body, and recommends appropriate question tags on Stack Overflow. Wang et al. developed an approach that combines topical interest, topical expertise, and activeness to recommend answers for new questions [47]. Similar to these studies, we investigate the potential of developing models for recommending information to highlight in SO answers. We adapted Named Entity Recognition (NER) models to identify the content that needs to be highlighted.
@@ 8 Conclusion
This paper is the first large-scale study of information highlighting in the text description of SO answers. We find that information highlighting is prevalent, with 47.6% of the 31,169,42 studied answers having text highlighted. We propose a terminology to categorize highlighted text in SO answers and find that source code content (e.g., identifiers and programming keywords) are frequently highlighted using Code and other highlighting formats (i.e., Bold and Italic). Besides code, users also tend to highlight updates (e.g., updates of answers), caveats (i.e., a reminder or warning of in which context or condition the provided solution works or does not work), and references. Further studies could put more effort into investigating how to use the highlighted content for downstream tasks (e.g., answer summarizing) that leverage information from SO answers, and provide tools that can suggest highlighted text for SO users. In addition, we also investigate the potential of recommending the content to be highlighted automatically by adopting named entity recognition (NER) models. Our experimental results highlight that our CNN-based models achieve precision scores between 0.5 and 0.72 across different formatting types, although they exhibit lower recall rates. On the other hand, while BERT demonstrates superior precision, it struggles with even lower recall. It’s worth noting that CNN-based Code model performs exceptionally well, with an impressive F1 score of 0.71, outperforming other models in our study. Although BERT usually has better performance in other tasks, in our case the CNN-based model worked better. In future research, we plan to investigate other more advanced models (e.g., LLaMA) and work on improving the recall rates for both models.